OpenStudy (anonymous):

The diameter of a construction pipe is 5ft . If the pipe is 20ft long, what is its volume?

OpenStudy (anonymous):

Volume = pi times radius squared times length

OpenStudy (anonymous):

@skullpatrol you should help her since you reported me giving the answers out to people when I am actually explaining to. But may I help this young lady? So YOU won't report me again ?

OpenStudy (anonymous):

pi(5^1)(20) = ?

OpenStudy (jdoe0001):

 \( \textit{volume of a cylinder} = \large \boldsymbol{\pi r^2 h}\)

OpenStudy (anonymous):

r=5/2 \[volume V=\pi r ^{2}h,\] h=20 plug the values and get the solution.